OpenStudy (anonymous):

Log[9](x+8)=log[9](x+7)=2

Directrix (directrix):

Log[9](x+8)=log[9](x+7)=2 ========= Is there supposed to be two = symbols in this problem?

OpenStudy (anonymous):

log[9](x+8) + log[9](x+7)=2

hero (hero):

Hint: log[9]((x+8)(x+7)) = 2

Directrix (directrix):

To log[9]((x+8)(x+7)) = 2, apply the definition of log. So, (x+8)(x+7) = ? @love4jazz

OpenStudy (anonymous):

(x+8)(x+7)=56

Directrix (directrix):

I agree with the left side but I did not get the 56. I'll check mine. Will you check how you got the 56 and report back? Thanks.

Directrix (directrix):

Using the log definition, I got 9 to the second power = (x+8)(x+7).